in jan. 2000 john was 11 times as old as his son. in jan 2012 he will be 3 times as old as his son. how old was his son in jan 2000?
Answer by ntnk(50)

Use S for the unknown age of the son:
In 2000:
Son: S
John: 11*S (11 times son's age)
12 Years Later:
Son: S + 12
John: 3(S+12) (3 times his sons age now) or (11*S)+12 (12 years older than in 2000)
3(S+12) is equal to 3*S + 3*12 because of the distributive property.
3*S + 36 = 11*S + 12
Subtract 12 from each side so it becomes:
3*S + 24 = 11*S
Now subtract 3*S from each side so it becomes:
24 = 8*S
Divide each side by 8:
3 = S
In 2000, the son is 3 and the father is 33.
In 2012, the son is 15 and the father is 45.
Check:
3*11=33
3+12=15
15*3=45
33+12=45

Solved By: NtNk